# Euclid Quick Data Release (Q1)

### v1

@@ 2.2 Euclidised lenses
In addition to the set of simulated images, we test the performance of LEMON using a sample of known strong gravitational lenses. These lenses are observed during several strong lensing surveys with HST or serendipitously discovered in the HST archive. Section 6.1 gives more details on the data set. We use the code HST2EUCLID  (Bergamini et al., in prep.) to Euclidise the HST images, mimicking Euclid observations of these sources at the same depth and resolution of the EWS in the $I_{\scriptscriptstyle\rm E}$ band.
@@
We summarise here the procedure implemented in HST2EUCLID. For more details, we refer the reader to Bergamini et al. (in prep.). First, we convert the pixel values of the HST images from units of electrons per second (${\rm e}\,{\rm s}^{-1}$) to physical flux densities (${\rm erg}\,{\rm s}^{-1}\,{\rm cm}^{-2}\,{\rm Hz}^{-1}$). We perform this operation for all available HST images in photometric bands overlapping with the $I_{\scriptscriptstyle\rm E}$ passband. For the lenses in the data set used in this paper, we use only HST observations with the Advanced Camera for Surveys (ACS) in the F814W band.
@@
In the second step, we account for the Euclid PSF. Since the input images are already convolved with the HST PSF, we do not convolve them directly with the Euclid PSF model. Instead, we create a convolution kernel that matches the HST to the Euclid PSF. Since we could not measure the PSF for each HST lens, we use a model for the ACS observations in the F814W band computed with the software TinyTim. [^2] https://www.stsci.edu/hst/instrumentation/focus-and-pointing/focus/tiny-tim-hst-psf-modeling The matched kernel for each pair of HST and Euclid PSFs is computed with the Python function create_matching_kernel of the Photutils package [^3] https://photutils.readthedocs.io/en/stable (Bradley et al. 2023). We then re-bin the images from the HST pixel grid to the correct Euclid pixel scale. The resulting images have pixel scales of $100\,{\rm mas}\,{\rm px}^{-1}$ in the $I_{\scriptscriptstyle\rm E}$ band. The re-binned images, expressed in units of physical flux densities, are then converted into units of electrons per second using the Euclid zero point.
@@
Finally, we add Poisson photon noise. To accomplish this task, we assume that the noise in the input HST background subtracted images is negligible compared to the noise in the Euclid observations. This approximation is supported by the significantly greater depth of the HST observations compared with the Euclid images. We adjust the noise level to achieve the nominal S/N of 10 for an extended source of magnitude 24.5 within an aperture of $0.65″$ radius expected for the EWS (Euclid Collaboration: Scaramella et al. 2022).
@@ 3 Introduction to LEMON
  • Aleatoric uncertainty: also known as statistical uncertainty, is associated with the intrinsic quality of the data analysed by the algorithm. The most common sources of aleatoric uncertainty are a low S/N, corruptions in the images (e.g., masked regions) and source blending. This kind of uncertainty cannot be reduced with training, because it depends on the quality of the image being analysed.
@@
The aleatoric uncertainty for a certain image is given as an output by LEMON, together with the predicted parameter value (for details on how the aleatoric uncertainty is predicted by LEMON, refer to Gentile et al. 2023). The epistemic uncertainty for an image can be obtained thanks to the dropout layers of LEMON: by randomly switching off some connections between the neurons of the neural network, predicting repeatedly the parameters associated to a single image is equivalent to sampling from the posterior probability of the parameters. The epistemic uncertainty can then be evaluated by measuring the standard deviation of this sampled distribution.
The total uncertainty for a generic parameter $p$ can be obtained via the following procedure (Gentile et al. 2023). For each image, LEMON predicts the mean value, $\overline{p}$, and the aleatoric uncertainty, $\sigma_{\textrm{A}}$, for $p$. A value $p_{i}$ is then sampled from a Gaussian distribution centred on $\overline{p}$ with a standard deviation equal to $\sigma_{\textrm{A}}$. This is repeated $\mathcal{N}$ times (we used $\mathcal{N}=50$ to optimise for speed), sampling the posterior distribution of the parameter $p$. Finally, we compute the standard deviation of the distribution associated with the extractions of $p_{i}$, obtaining the total uncertainty.
LEMON employs the ResNet architecture (He et al. 2015), a deep residual network known for its success on computer vision tasks (Russakovsky et al. 2015). The architecture of LEMON is composed of: convolutional layers, used for feature extraction; dropout layers, to facilitate Bayesian uncertainty estimation through Monte Carlo sampling; finally, a fully connected layer made by $2N_{\textrm{param}}$ nodes, with $N_{\textrm{param}}$ number of parameters estimated by the network. A corresponding number of aleatoric uncertainties is estimated through the minimisation of a Gaussian log-likelihood loss function, with an additional regularisation term in order to prevent overly large or small variance estimates. Technical details on the implementation can be found in Gentile et al. (2023).
@@ 4 Methods
In this section, we will first describe the training procedure for LEMON in Sect. 4.1. Section 4.2 is dedicated to the description of the metrics that we used to evaluate the performance of LEMON. Finally, Sect. 4.3 is dedicated to discussing the calibration procedure for the total uncertainty.
@@ 4.1 Training the network
We prepare the training of LEMON by taking the full sample of 100 000100000100\,000100 000 mock Euclid lenses, both with and without companions in the image, and splitting it into a training set formed by 80 0008000080\,00080 000 lenses (equally split between companions and non-companions), a validation set and a test set formed by 10 0001000010\,00010 000 lenses each (equally split as before). We shuffle the three sets to avoid possible biases during learning. Differently from Gentile et al. (2023), we use the ResNet-50 architecture (He et al. 2015) when implementing LEMON. The increased number of layers with respect to ResNet-34 should improve the accuracy in the recovery of the parameters. Before feeding the images into LEMON, we apply a preprocessing step where pixel values are first rescaled to the range [0,1], and then transformed using a square root scaling to enhance the visibility of low-intensity features.
Training is performed via stochastic gradient descent, with a batch size of 64 images. For the optimiser, we used ADAM (Kingma & Ba 2015), with a starting learning rate of $10^{-4}$. To prevent overfitting, we used both the ReduceLROnPlateau and EarlyStopping callbacks of the keras Python library. The former reduces the learning rate of the optimiser when no improvement is seen for a given number of epochs, while the latter stops the training when the validation loss stops decreasing. Dropout layers, needed for estimating the epistemic uncertainty, are employed at both training and testing time, with a drop rate of 0.05.
We performed the training via the tensorflow Python library (Abadi et al. 2015), using a single Nvidia RTX 4080 graphics processing unit (GPU). The whole training procedure required 15 hours and 30 minutes, with a total of 70 epochs.
@@ 4.2 Metrics
To evaluate the performance of LEMON in recovering the lens parameters, we make use of various statistical estimators.
  • Bias, defined as
    $\mu(\hat{\mathbf{y}},\mathbf{y}):=\textrm{median}(\hat{\mathbf{y}}-\mathbf{y})\,,$ (4)
    where $\hat{\mathbf{y}}$ and $\mathbf{y}$ are the arrays of the predicted and the true parameter values, respectively. [^4] A single predicted value $\hat{y}_{i}$ is the mean of the distribution of the $\mathcal{N}$ parameter extractions $p_{j}$ described in Sect. 3. This is an indicator of the accuracy of the prediction: higher bias values indicate a systematic overestimation of the predictions, and vice versa for lower bias values.
  • Discrepancy distribution scatter: the scatter of the discrepancy distribution is obtained from the 16th and 84th percentiles of the discrepancy distribution via the following definitions:
    $\sigma^{-}$ $:=\textrm{median}(\hat{\mathbf{y}}-\mathbf{y})-\textrm{percentile}_{\textrm{16th}}(\hat{\mathbf{y}}-\mathbf{y})\,,$ (5)
    $\sigma^{+}$ $:=\textrm{percentile}_{\textrm{84th}}(\hat{\mathbf{y}}-\mathbf{y})-\textrm{median}(\hat{\mathbf{y}}-\mathbf{y})\,.$ (6)
    Bigger values of these metrics indicate higher scatter from the ideal $\hat{y}=y$ relation. Since they are obtained from the percentiles, these quantities are robust estimators of the scatter.
@@
    where $\hat{y}_{i}$ and $y_{i}$ are the generic components of the $\hat{\mathbf{y}}$ and $\mathbf{y}$ arrays, respectively, which have length $N$. This metric is a measure of predictive power: a value of zero indicates a perfect fit to the true parameter values.
  • Mean absolute error (MAE), defined as
    $\textrm{MAE}:=\frac{1}{N}\sum_{i=1}^{N}\left|\hat{y}_{i}-y_{i}\right|\,.$ (8)
    This metric is similar, but not equivalent to RMSE. It is the average absolute vertical distance between each point $(y_{i},\hat{y}_{i})$ and the ideal $\hat{y}=y$ relation.
@@
    $\textrm{NMAD}:=1.48\,\textrm{median}[\left|(\hat{\mathbf{y}}-\mathbf{y})-\textrm{median}(\hat{\mathbf{y}}-\mathbf{y})\right|]\,.$ (9)
    Figure 3: Reliability plots for the parameters predicted by LEMON. Left: trends before the calibration procedure. Right: trends after the calibration procedure.
    The NMAD is roughly analogous to the standard deviation ($\sigma\approx\kappa\,\textrm{MAD}$, where MAD is the median absolute deviation and $\kappa$ is a constant scale factor, which is approximately equal to $1.48$ for normally distributed data). An advantage of the NMAD metric is that it is a non-parametric estimator, which is robust to outliers.
@@
    is the mean value of the true parameter over all the lenses. This metric provides a measure of how well the parameters are predicted by LEMON. Given a fixed parameter, if it is perfectly predicted by the network, $\hat{y}_{i}=y_{i}$ for all the points of the data set, and thus $R^{2}=1$. A baseline network, which always outputs the mean value of the true parameter (i.e., $\hat{y}_{i}=\overline{y}$), will produce $R^{2}=0$. If the network performs worse than the baseline model (i.e., the mean of the data provides a better fit than than the predicted values), the values of $R^{2}$ will be negative. [^5] Notice that the coefficient of determination is not the square of a quantity $R$, and therefore can also assume negative values.
@@ 4.3 Calibration procedure
Uncertainties produced by BNNs are known to be affected by a systematic over- or under-estimation of the confidence intervals (Guo et al. 2017). Therefore, a calibration procedure for the uncertainties is required. Following Gentile et al. (2023), we implement the Platt-scaling method (Kull et al. 2017), which consists in rescaling the predicted total uncertainty of a given parameter by a factor $s:\sigma\rightarrow s\sigma$, in order to match the empirical cumulative distribution function (CDF) with the CDF of a Gaussian distribution.
Figure 4: Recovery plot of the parameters of Euclid mock lenses, showing how well the predicted model parameters from LEMON ($y$ coordinate) reproduce the corresponding true value ($x$ coordinate), for each lens of the test set (blue points). The ideal recovery line is shown as a dashed black line. The median trend of the scatter plots, along with the respective scatters associated with the 16th and 84th percentiles, are shown as red points and bars, respectively. For each panel, we also report the respective cumulative metrics. The bottom two rows of each parameter show the bias and the NMAD, respectively, as defined in Sect. 4.2, as a function of the true value of the parameters estimated by LEMON. For clarity, we have plotted only $2000$ random points from the test set.
To obtain the scaling factors, we first proceed by splitting the validation set into $70\%$ for the calibration training set and $30\%$ for the calibration test set. We then build the uncalibrated reliability plot, by putting on the $x$-axis the values extracted from the CDF of a Gaussian distribution and on the $y$-axis the respective values extracted from the actual CDF. We fit the resulting trend with a $\beta$-function, and find the Platt-scaling factor, $s$, by minimising the difference between the $\beta$-function and the bisector of the reliability plot. This is done for each parameter, obtaining the values reported in Table 2.
@@
Einstein radius $R_{\textrm{Ein}}/\textrm{arcsec}$ $0.99$
Axis ratio $q$ $0.97$
[…]
Lens effective radius $R_{\textrm{e, lens}}/\textrm{arcsec}$ $1.01$
Lens Sérsic index $n_{\textrm{lens}}$ $1.07$
Lens magnitude $m_{\textrm{lens}}$ $1.02$
Figure 3 shows the uncalibrated reliability plot for each parameter in the left panel, and the calibrated reliability plot on the right panel. We can see from the figure that the calibration procedure has its biggest effect on the position angle and Sérsic index uncertainties, which were previously systematically overestimated and underestimated, respectively.

### v2

@@ 2.2 Euclidised lenses
In addition to the set of simulated images, we tested the performance of LEMON using a sample of known strong gravitational lenses. These lenses were observed during several strong lensing surveys with HST or serendipitously discovered in the HST archive. Section 6.1 gives more details on the dataset. We used the code HST2EUCLID (Euclid Collaboration: Bergamini et al. 2025) to Euclidise the HST images, mimicking Euclid observations of these sources at the same depth and resolution of the EWS in the $I_{\scriptscriptstyle\rm E}$ band.
@@
We summarise here the procedure implemented in HST2EUCLID. For more details, we refer the reader to Euclid Collaboration: Bergamini et al. (2025). First, we converted the pixel values of the HST images from units of electrons per second (${\rm e}\,{\rm s}^{-1}$) to physical flux densities (${\rm erg}\,{\rm s}^{-1}\,{\rm cm}^{-2}\,{\rm Hz}^{-1}$). We performed this operation for all available HST images in photometric bands overlapping with the $I_{\scriptscriptstyle\rm E}$ passband. For the lenses in the dataset used in this paper, we used only HST observations with the Advanced Camera for Surveys (ACS) in the F814W band. Notice that the F814W band does not perfectly overlap with the $I_{\scriptscriptstyle\rm E}$ band. The Euclidisation procedure, however, assumes a constant spectral energy density (SED) for the lens galaxy, such that the magnitude in the $I_{\scriptscriptstyle\rm E}$ band is equal to the one in the F814W band. This means that there is no photon loss, even if the F814W band is narrower than the $I_{\scriptscriptstyle\rm E}$ band.
@@
In the second step, we accounted for the Euclid PSF. The Euclid PSF used by HST2EUCLID in this procedure coincides with the one used to produce the Euclid mock lenses described in Sect. 2.1, and was derived during the pre-launch VIS and NISP detector characterisation. This model matches well the true Euclid PSFs. For more details on the PSF, we refer the reader to Euclid Collaboration: Bergamini et al. (2025). Since the input images were already convolved with the HST PSF, we did not convolve them directly with the Euclid PSF model. Instead, we created a convolution kernel that matches the HST to the Euclid PSF. Since we could not measure the PSF for each HST lens, we used a model for the ACS observations in the F814W band computed with the software TinyTim. [^2] https://www.stsci.edu/hst/instrumentation/focus-and-pointing/focus/tiny-tim-hst-psf-modeling The matched kernel for each pair of HST and Euclid PSFs was computed with the Python function create_matching_kernel of the Photutils package [^3] https://photutils.readthedocs.io/en/stable (Bradley et al. 2023). We then re-binned the images from the HST pixel grid to the correct Euclid pixel scale. The resulting images have pixel scales of $100\,{\rm mas}\,{\rm px}^{-1}$ in the $I_{\scriptscriptstyle\rm E}$ band. The re-binned images, expressed in units of physical flux densities, were then converted into units of electrons per second using the Euclid zero point.
@@
Finally, we added Poisson photon noise. To accomplish this task, we assumed that the noise in the input HST background subtracted images is negligible compared to the noise in the Euclid observations. This approximation is supported by the significantly greater depth of the HST observations compared with the Euclid images. We adjusted the noise level to achieve the nominal S/N of 10 for an extended source of magnitude 24.5 within an aperture of 0.65 radius expected for the EWS (Euclid Collaboration: Scaramella et al. 2022).
@@ 3 Introduction to LEMON
  • Aleatoric uncertainty: also known as statistical uncertainty, is associated with the intrinsic quality of the data analysed by the algorithm. The most common sources of aleatoric uncertainty are a low S/N, corruptions in the images (e.g. masked regions), and source blending. This kind of uncertainty cannot be reduced with training, because it depends on the quality of the image being analysed.
@@
The aleatoric uncertainty for a certain image is given as an output by LEMON, together with the predicted parameter value (for details on how the aleatoric uncertainty is predicted by LEMON, refer to Gentile et al. 2023). The epistemic uncertainty for an image can be obtained thanks to the dropout layers of LEMON: by randomly switching off some connections between the neurons of the neural network, predicting repeatedly the parameters associated with a single image is equivalent to sampling from the posterior probability of the parameters. The epistemic uncertainty can then be evaluated by measuring the standard deviation of this sampled distribution.
The total uncertainty for a generic parameter, $p$, can be obtained via the following procedure (Gentile et al. 2023). For each image, LEMON predicts the mean value, $\overline{p}$, and the aleatoric uncertainty, $\sigma_{\textrm{A}}$, for $p$. A value $p_{i}$ is then sampled from a Gaussian distribution centred on $\overline{p}$ with a standard deviation equal to $\sigma_{\textrm{A}}$. This is repeated $\mathcal{N}$ times (we used $\mathcal{N}=50$ to optimise for speed), sampling the posterior distribution of the parameter $p$. Finally, we compute the standard deviation of the distribution associated with the extractions of $p_{i}$, obtaining the total uncertainty.
LEMON employs the ResNet architecture (He et al. 2016), a deep residual network known for its success on computer vision tasks (Russakovsky et al. 2015). The architecture of LEMON is composed of: convolutional layers, used for feature extraction; dropout layers, to facilitate Bayesian uncertainty estimation through Monte Carlo sampling; finally, a fully connected layer made by $2N_{\textrm{param}}$ nodes, with $N_{\textrm{param}}$ number of parameters estimated by the network. A corresponding number of aleatoric uncertainties is estimated through the minimisation of the following Gaussian log-likelihood loss function:
$\ln\mathcal{L}(\textbf{y},\hat{\textbf{y}})=\sum_{j=1}^{N_{\textrm{P}}}\left[-\frac{1}{2\sigma^{2}_{i,j}}\lVert y_{i,j}-\hat{y}_{i,j}\rVert-\frac{1}{2}\ln(\sigma^{2}_{i,j})\right]\,$, (6)
where $\hat{\mathbf{y}}$ and $\mathbf{y}$ are the arrays of the predicted and the true parameter values, respectively, [^4] A single predicted value $\hat{y}_{i}$ is the mean of the distribution of the $\mathcal{N}$ parameter extractions $p_{j}$. while $\sigma_{i,j}$ is the aleatoric uncertainty predicted for the parameter $j$ of image $i$. The second term in Eq. (6) is an additional regularisation term, introduced to prevent overly large or small variance estimates. Additional details on the implementation can be found in Gentile et al. (2023).
@@ 4 Methods
In this section we first describe the training procedure for LEMON in Sect. 4.1. Section 4.2 is dedicated to the description of the metrics that we used to evaluate the performance of LEMON. Finally, Sect. 4.3 is dedicated to discussing the calibration procedure for the total uncertainty.
@@ 4.1 Training the network
We prepared the training of LEMON by taking the full sample of 100 000100\,000 mock Euclid lenses and splitting it into a training set formed of 80 00080\,000 lenses, a validation set, and a test set formed by 10 00010\,000 lenses each. We shuffled the three sets to avoid possible biases during learning. Unlike Gentile et al. (2023), we used the ResNet-50 architecture (He et al. 2016) when implementing LEMON. We verified that this choice does not influence the results, confirming the analysis of Wagner-Carena et al. (2024), which shows that marginal improvements over the ResNet-34 architecture happen only at very large training set sizes. Before feeding the images into LEMON, we applied a pre-processing step whereby pixel values were first rescaled to the range [0,1], and then transformed using a square root scaling to enhance the visibility of low-intensity features.
Training was performed via stochastic gradient descent, with a batch size of 64 images. For the optimiser, we used ADAM (Kingma & Ba 2015), with a starting learning rate of $10^{-4}$. To prevent overfitting, we used both the ReduceLROnPlateau and EarlyStopping callbacks of the keras Python library. The former reduces the learning rate of the optimiser when no improvement is seen for a given number of epochs, while the latter stops the training when the validation loss stops decreasing. Dropout layers, needed for estimating the epistemic uncertainty, are employed at both training and testing time, with a drop rate of 0.05. We performed the training via the tensorflow Python library (Abadi et al. 2015), using a single Nvidia RTX 4080 graphics processing unit (GPU).
@@ 4.2 Metrics
To evaluate the performance of LEMON in recovering the lens parameters, we made use of various statistical estimators.
  • Bias, defined as [^5] The writing $\textrm{median}\{x_{i}|i\}$ denotes the median of the set $\{x_{i}|i=1,\dots,N\}$ of vector components $x_{i}$.
    $\mu:=\textrm{median}\{\hat{y}_{i}-y_{i}|i\}\,.$ (7)
    This is an indicator of the accuracy of the prediction: higher bias values indicate a systematic overestimation of the predictions, and vice versa for lower bias values. For comparisons with forward modelling, we implement the uncertainty on the estimates by using the weighted version of the median, with weights equal to the reciprocal of the variance associated with the predictions.
@@
    where $\hat{y}_{i}$ and $y_{i}$ are the generic components of the $\hat{\mathbf{y}}$ and $\mathbf{y}$ arrays, respectively, which have length $N$. This metric is a measure of predictive power: a value of zero indicates a perfect fit to the true parameter values. For comparisons with forward modelling, we implement the uncertainty on the estimates by using the function root_mean_squared_error from the scikit-learn Python library (which natively supports weights for the inputs), with weights equal to the reciprocal of the variance associated with the predictions.
@@
    $\textrm{NMAD}:=1.48\,\textrm{median}\{\left|(\hat{y}_{i}-y_{i})-\textrm{median}\{\hat{y}_{i}-y_{i}|i\}\right|\,|i\}\,.$ (9)
    Figure 3: Reliability plots for the parameters predicted by LEMON. The lines represent, for each parameter, the discrepancy between the ideal cumulative statistical coverage from a Gaussian distribution ($x$ axis) and the corresponding empirical one ($y$ axis). The closer the curves are to the ideal relation (dashed black line), the more reliable the uncertainty estimates are. Left: Trends before the calibration procedure. Right: Trends after the calibration procedure.
    The NMAD is roughly analogous to the standard deviation ($\sigma\approx\kappa\,\textrm{MAD}$, where MAD is the median absolute deviation and $\kappa$ is a constant scale factor, which is approximately equal to $1.48$ for normally distributed data). For comparisons with forward modelling, we substitute the median with the weighted median, with weights equal to the reciprocal of the variance associated with the predictions.
@@
    is the mean value of the true parameter over all the lenses. This metric provides a measure of how well the parameters are predicted by LEMON. Given a fixed parameter, if it is perfectly predicted by the network, $\hat{y}_{i}=y_{i}$ for all the points of the dataset, and thus $R^{2}=1$. A baseline network, which always outputs the mean value of the true parameter (i.e. $\hat{y}_{i}=\overline{y}$), will produce $R^{2}=0$. If the network performs worse than the baseline model (i.e. the mean of the data provides a better fit than than the predicted values), the values of $R^{2}$ will be negative. [^6] Notice that the coefficient of determination is not the square of a quantity $R$, and therefore can also assume negative values. For comparisons with forward modelling, we implement the uncertainty on the estimates by using the function r2_score from the scikit-learn Python library (which natively supports weights for the inputs), with weights equal to the reciprocal of the variance associated with the predictions.
@@ 4.3 Calibration procedure
Uncertainties produced by BNNs are known to be affected by a systematic over- or under-estimation of the confidence intervals (Guo et al. 2017). Therefore, a calibration procedure for the uncertainties is required. Following Gentile et al. (2023), we implement the Platt-scaling method (Kull et al. 2017), which consists of rescaling the predicted total uncertainty of a given parameter by a factor $s:\sigma\rightarrow s\sigma$, in order to match the empirical cumulative distribution function (CDF) with the CDF of a Gaussian distribution.
Figure 4: Recovery plot of the parameters of Euclid mock lenses, showing how well the predicted model parameters from LEMON ($y$ co-ordinate) reproduce the corresponding true value ($x$ co-ordinate), for each lens of the test set (blue points). The ideal recovery line is shown as a dashed black line. The median trend of the scatter plots, along with the respective scatters associated with the 16th and 84th percentiles, are shown as red points and bars, respectively. For each panel, we also report the respective cumulative metrics. The bottom two rows of each parameter show the bias and the NMAD, respectively, as defined in Sect. 4.2, as a function of the true value of the parameters estimated by LEMON. For clarity, we have plotted only $5000$ random points from the test set.
To obtain the scaling factors, we first proceeded by splitting the validation set into $70\%$ for the calibration training set and $30\%$ for the calibration test set. We then built the uncalibrated reliability plot, by putting on the $x$ axis the values extracted from the CDF of a Gaussian distribution and on the $y$ axis the respective values extracted from the actual CDF. We fitted the resulting trend with a $\beta$-function, and found the Platt-scaling factor, $s$, by minimising the difference between the $\beta$-function and the bisector of the reliability plot. This was done for each parameter, obtaining the values reported in Table 2. Figure 3 shows the uncalibrated reliability plot for each parameter in the left panel, and the calibrated reliability plot on the right panel.
It should be noted that the Platt-scaling calibration we use for the uncertainties has been performed on a dataset of simulated Euclid-like lenses. As such, while this correction may hold for similar lenses, it could be not robust when switching to real data. An analysis of the robustness of this method on real lenses is discussed in Appendix C.
@@
Einstein radius $R_{\textrm{Ein}}/\textrm{arcsec}$ $0.84$
Mass $x$-ellipticity $\epsilon_{x,\mathrm{m}}$ $0.92$
Mass $y$-ellipticity $\epsilon_{y,\mathrm{m}}$ $0.91$
Light $x$-ellipticity $\epsilon_{x,\mathrm{l}}$ $0.78$
Light $y$-ellipticity $\epsilon_{y,\mathrm{l}}$ $0.79$
Lens effective radius $R_{\textrm{e, lens}}/\textrm{arcsec}$ $0.89$
Lens Srsic index $n_{\textrm{lens}}$ $0.85$
Lens magnitude $m_{\textrm{lens}}$ $0.79$
Source $x$ position $x_{\textrm{src}}/\textrm{arcsec}$ $0.88$
Source $y$ position $y_{\textrm{src}}/\textrm{arcsec}$ $0.88$